Ryza borrows P 150,000. At what rate compounded bi – monthly will her interest be if she agrees to pay P10,000 more at the end of 2 years?



1
Expert's answer
2021-12-02T16:16:07-0500

"FV=PV(1+r/6)^{2\cdot6}"

"(150000+10000)=150000(1+r/6)^{2\cdot6}"

"r=6((16/15)^{1/12}-1)=0.032=3.2\%"